# Collaborative Agents for Software Engineering

### v2

@@ Appendix A Complete Related Work
Automating Code Review Activities Our focus included detecting source code vulnerabilities, ensuring style alignment, and maintaining commit message and code consistency. Other studies explore various aspects of code review. Hellendoorn et al. [15] addressed the challenge of anticipating code change positions. Siow et al. [29] introduced CORE, employing multi-level embeddings for code modification semantics and retrieval-based review suggestions. Hong et al. [18] proposed COMMENTFINDER, a retrieval-based method for suggesting comments during code reviews. Tufano et al. [35] designed T5CR with SentencePiece, enabling work with raw source code without abstraction. Li et al. [23] developed CodeReviewer, focusing on code diff quality, review comment generation, and code refinement using the T5 model. Recently, large language models have been incorporated; Lu et al. [25] fine-tuned LLama with prefix tuning for LLaMA-Reviewer, using parameter-efficient fine-tuning and instruction tuning in a code-centric domain.
@@
Collaborative AI Collaborative AI refers to artificial intelligent systems designed to achieve shared goals with humans or other AI systems. Previous research extensively explores the use of multiple LLMs in collaborative settings, as demonstrated by Talebirad et al. [31] and Qian et al. [28]. These approaches rely on the idea that inter-agent interactions enable LLMs to collectively enhance their capabilities, leading to improved overall performance. The research covers various aspects of multi-agent scenarios, including collective thinking, conversation dataset curation, sociological phenomenon exploration, and collaboration for efficiency. Collective thinking aims to boost problem-solving abilities by orchestrating discussions among multiple agents. Researchers like Wei et al. [39] and Li et al. [21] have created conversational datasets through role-playing methodologies. Sociological phenomenon investigations, such as Park et al. [27]’s work, involve creating virtual communities with rudimentary language interactions and limited cooperative endeavors. In contrast, Akata et al. [1] scrutinized LLM cooperation through orchestrated repeated games. Collaboration for efficiency, proposed by Cai et al. [5], introduces a model for cost reduction through large models as tool-makers and small models as tool-users. Zhang et al. [42] established a framework for verbal communication and collaboration, enhancing overall efficiency. However, Li et al. [21] and Qian et al. [28], presenting a multi-agent framework for software development, primarily relied on natural language conversations, not standardized software engineering documentation, and lacked advanced human process management expertise. Challenges in multi-agent cooperation include maintaining coherence, avoiding unproductive loops, and fostering beneficial interactions. Our approach emphasizes integrating advanced human processes, like code review in software maintenance, within multi-agent systems.
Appendix B Experimental Details
@@ B.1 Role Definition
B.2 Execute Time Across Languages
@@
Appendix C Comparative Analysis of QA-Checker AI System and Recursive Self-Improvement Systems
In this section, we will delve into the differences between QA-Checker and self-improvement systems [17], and underscore the importance of the QA-Checker in role conversations.
C.1 Comparison Table
@@
| Feature/System | QA-Checker AI System | Recursive Self-Improvement System |
@@
Specialized for QA tasks with precise task execution |
Broad scope, covering various dimensions like software development and learning algorithms |
@@
Advanced optimization techniques for iterative improvement in QA |
Multi-level learning: learning, meta-learning, and recursive self-improvement |
@@
Focused on individual capability in specific QA tasks |
Enhances the entire system, including multi-agent interactions and communication protocols |
@@
Based on mathematical models to optimize answer quality |
Utilizes experiences from past projects to improve overall performance |
C.2 Differences and Implications
@@
C.3 Importance of QA-Checker in Role Conversations
@@ C.4 Conclusion
Appendix D Capabilities Analysis between CodeAgent and Other Methods
Compared to open-source baseline methods such as AutoGPT and autonomous agents such as ChatDev and MetaGPT, CodeAgent offers functions for code review tasks: consistency analysis, vulnerability analysis, and format analysis. As shown in Table 8, our CodeAgent encompasses a wide range of abilities to handle complex code review tasks efficiently. Incorporating the QA-Checker self-improved module can significantly improve the conversation generation between agents and contribute to the improvement of code review. Compared to COT, the difference and the advantages of CodeAgent with QA-Checker are shown in Section C.
Appendix E Dataset
@@
As shown in [44], our study incorporates three distinct datasets for evaluating the performance of CodeAgent: , , and . , compiled by Tufano et al. [35], derives from Gerrit and GitHub projects, excluding noisy or overly lengthy comments and review data with new tokens in revised code not present in the initial submission. , collected by Thongtanunam et al. [32] from three Gerrit repositories, comprises only submitted and revised codes without review comments. Lastly, , gathered by Tufano et al. [34] from Java projects on GitHub, filters out noisy, non-English, and duplicate comments. These datasets are employed for Code Revision Before Review (CRB) and Code Revision After Review (CRA) tasks, with the exception of for CRA and Review Comment Generation (RCG) due to its lack of review comments.
@@
Our dataset, illustrated in Fig. 8, encapsulates a detailed analysis of consistency and format detection in software development, spanning various programming languages. It includes CA (consistency between commit and commit message (See Table 1)) and FA (format consistency between commit and original (See Table 1)) data, segmented into positive and negative samples based on the merged and closed status of pull requests. For example, in Python, the dataset comprises 254 merged and 35 closed negative CA samples, alongside 803 merged and 213 closed positive CA samples, with corresponding distributions for other languages like Java, Go, C++, and more. Similarly, the FA data follows this pattern of positive and negative samples across languages. Figure 7 graphically represents this data, highlighting the distribution and comparison of merged versus closed samples in both CA and FA categories for each language. This comprehensive dataset, covering over 3,545 commits and nearly 2,933 pull requests from more than 180 projects, was meticulously compiled using a custom crawler designed for GitHub API interactions, targeting post-April 2023 repositories to ensure up-to-date and diverse data for an in-depth analysis of current software development trends.
@@
Appendix F Key Factors Leading to Vulnerabilities
@@
Check for vulnerabilities like SQL injection, Cross-Site Scripting (XSS), and command injection in new or modified code, especially where user input is processed. |
| 2 |
Buffer Overflows |
@@
Particularly in lower-level languages, ensure that memory management is handled securely to prevent overflows. |
| 3 |
Authentication and Authorization Flaws |
@@
Evaluate any changes in authentication and authorization logic for potential weaknesses that could allow unauthorized access or privilege escalation. |
| 4 |
Sensitive Data Exposure |
@@
Assess handling and storage of sensitive information like passwords, private keys, or personal data to prevent exposure. |
| 5 |
Improper Error and Exception Handling |
@@
Ensure that errors and exceptions are handled appropriately without revealing sensitive information or causing service disruption. |
| 6 |
Vulnerabilities in Dependency Libraries or Components |
@@
Review updates or changes in third-party libraries or components for known vulnerabilities. |
| 7 |
Cross-Site Request Forgery (CSRF) |
@@
Verify that adequate protection mechanisms are in place against CSRF attacks. |
| 8 |
Unsafe Use of APIs |
@@
Check for the use of insecure encryption algorithms or other risky API practices. |
| 9 |
Code Injection |
Look for vulnerabilities related to dynamic code execution. |
| 10 |
Configuration Errors |
@@
Ensure that no insecure configurations or settings like open debug ports or default passwords have been introduced. |
| 11 |
Race Conditions |
@@
Analyze for potential data corruption or security issues arising from race conditions. |
| 12 |
Memory Leaks |
@@
Identify any changes that could potentially lead to memory leaks and resource exhaustion. |
| 13 |
Improper Resource Management |
@@
Check resource management, such as proper closure of file handles or database connections. |
| 14 |
Inadequate Security Configurations |
@@
Assess for any insecure default settings or unencrypted communications. |
| 15 |
Path Traversal and File Inclusion Vulnerabilities |
@@
Examine for risks that could allow unauthorized file access or execution. |
| 16 |
Unsafe Deserialization |
@@
Look for issues that could allow the execution of malicious code or tampering with application logic. |
| 17 |
XML External Entity (XXE) Attacks |
Check if XML processing is secure against XXE attacks. |
| 18 |
Inconsistent Error Handling |
@@
Review error messages to ensure they do not leak sensitive system details. |
| 19 |
Server-Side Request Forgery (SSRF) |
@@
Analyze for vulnerabilities that could be exploited to attack internal systems. |
| 20 |
Unsafe Redirects and Forwards |
@@
Check for vulnerabilities leading to phishing or redirection attacks. |
| 21 |
Use of Deprecated or Unsafe Functions and Commands |
@@
Identify usage of any such functions and commands in the code. |
| 22 |
Code Leakages and Hardcoded Sensitive Information |
@@
Look for hardcoded passwords, keys, or other sensitive data in the code. |
| 23 |
Unencrypted Communications |
@@
Verify that data transmissions are securely encrypted to prevent interception and tampering. |
| 24 |
Mobile Code Security Issues |
@@
For mobile applications, ensure proper handling of permission requests and secure data storage. |
| 25 |
Cloud Service Configuration Errors |
@@
Review any cloud-based configurations for potential data leaks or unauthorized access. |
Appendix G Data Leakage Statement
@@
As the new dataset introduced in Section E, the time of the collected dataset is after April 2023, avoiding data leakage while we evaluate CodeAgent on codeData dataset.
Appendix H Algorithmic Description of CodeAgent Pipeline with QA-Checker
@@
Appendix I Detailed Performance of CodeAgent in Various Languages on VA task
In our comprehensive analysis using CodeAgent, as detailed in Table 11, we observe a diverse landscape of confirmed vulnerabilities across different programming languages. The table categorizes these vulnerabilities into ‘merged’ and ‘closed’ statuses for languages such as Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. A significant finding is a markedly high number of ‘merged’ vulnerabilities in Python, potentially reflective of its extensive application or intrinsic complexities leading to security gaps. Conversely, languages like Go, Ruby, and C exhibit notably lower counts in both categories, perhaps indicating lesser engagement in complex applications or more robust security protocols. Table 11 that the ‘closed’ category consistently presents lower vulnerabilities than ‘merged’ across most languages, signifying effective resolution mechanisms. However, an exception is noted in C, where ‘closed’ counts surpass those of ‘merged’, possibly indicating either delayed vulnerability identification or efficient mitigation strategies. Remarkably, the Rate is generally observed to be higher than Rate across the languages, exemplifying a significant reduction in vulnerabilities post-resolution. For example, Python demonstrates a Rate of 14.00% against a higher Rate of 18.16%. This trend is consistent in most languages, emphasizing the importance of proactive vulnerability management. The Rate, representing the proportion of confirmed vulnerabilities against the total of both merged and closed items, further elucidates this point, with C++ showing the highest Rate at 16.49%. These insights not only underline the diverse vulnerability landscape across programming languages but also highlight the adeptness of CodeAgent in pinpointing and verifying vulnerabilities in these varied contexts.
@@ Appendix I Detailed Performance of CodeAgent in Various Languages on VA task
| Rate | 14.79% | 7.03% | 7.73% | 16.49% | 12.76% | 13.46% | 10.45% | 14.47% | 9.73% |
Appendix J More detailed experimental results on CA and FA tasks
@@
Appendix K Case Study
K.1 Performance on 9 languages
K.2 Difference of CodeAgent-3.5 and CodeAgent-4.0
CodeAgent-3.5 and CodeAgent-4.0 in this paper has no difference in general code review, however, as CodeAgent-4.0 is more powerful in processing long input sequences and logic reasoning. As shown in Figure 22, we take one example of consistency detection between commit and commit message and find that CodeAgent-4.0 diffs from CodeAgent-3.5 in the detailed explanation. CodeAgent-3.5 output a report with 15k lines while CodeAgent-4.0 outputs a report with more than 17.7k lines. Detailed data is shown in https://zenodo.org/records/10607925.
Appendix L Ablation study
In this section, we evaluate the performance of different parts in CodeAgent in vulnerability analysis. CodeAgent is based on chain-of-thought (COT) and large language model (a.k.a. ChatGPT). As shown in Section 5.1, CodeAgent outperforms baselines (a.k.a. CodeBERT, ChatGPT-3.5, ChatGPT-4.0) across 9 different languages. The performance mainly comes from the combination of COT and QA-Checker. Thus, we design an additional version called CodeAgent , which means CodeAgent without QA-Checker. Then, we use CodeAgent to do vulnerability analysis and compare with CodeAgent. We first discuss about the result of CodeAgent and then discuss about comparison between CodeAgent and CodeAgent .
Overview of Vulnerabilities in CodeAgent
Table 12 presents the findings of CodeAgent , a variant of the original CodeAgent, in identifying vulnerabilities across different programming languages. The table showcases the number of ‘merged’ and ‘closed’ vulnerabilities in languages such as Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. Notably, Python leads in the ‘merged’ category with a total of 1,057 cases, of which 140 are confirmed, yielding a Rate of 13.25%. In contrast, languages like Go and Ruby show lower vulnerability counts in both ‘merged’ and ‘closed’ categories. The table also includes Rate and Rate, providing insights into the effectiveness of vulnerability management across these languages.
Detailed Comparison between CodeAgent and CodeAgent
Comparing the findings in Table 12 with those in Table 11, we observe some notable differences in vulnerability detection by CodeAgent and CodeAgent . While the overall trend of higher ‘merged’ vulnerabilities in Python and lower counts in Go and Ruby remains consistent, Table 12 shows a slight reduction in the Rate for most languages, suggesting a more conservative confirmation approach in CodeAgent . Similarly, Rate and Rate values in Table 12 generally indicate a lower proportion of confirmed vulnerabilities compared to Table 11, reflecting potentially different criteria or efficacy in vulnerability assessment. These variations highlight the impact of QA-Checker in CodeAgent.
@@
| Rate | 13.49% | 6.77% | 7.25% | 12.37% | 11.22% | 13.46% | 10.45% | 15.47% | 8.56% |
Appendix M Tool

### v4

@@ Appendix B Complete Related Work
Automating Code Review Activities Our focus included detecting source code vulnerabilities, ensuring style alignment, and maintaining commit message and code consistency. Other studies explore various aspects of code review. Hellendoorn et al. Hellendoorn et al. (2021) addressed the challenge of anticipating code change positions. Siow et al. Siow et al. (2020) introduced CORE, employing multi-level embeddings for code modification semantics and retrieval-based review suggestions. Hong et al. Hong et al. (2022) proposed COMMENTFINDER, a retrieval-based method for suggesting comments during code reviews. Tufano et al. Tufano et al. (2021) designed T5CR with SentencePiece, enabling work with raw source code without abstraction. Li et al. Li et al. (2022) developed CodeReviewer, focusing on code diff quality, review comment generation, and code refinement using the T5 model. Recently, large language models have been incorporated; Lu et al. Lu et al. (2023) fine-tuned LLama with prefix tuning for LLaMA-Reviewer, using parameter-efficient fine-tuning and instruction tuning in a code-centric domain.
@@
Collaborative AI Collaborative AI refers to artificial intelligent systems designed to achieve shared goals with humans or other AI systems. Previous research extensively explores the use of multiple LLMs in collaborative settings, as demonstrated by Talebirad et al. Talebirad and Nadiri (2023) and Qian et al. Qian et al. (2023). These approaches rely on the idea that inter-agent interactions enable LLMs to collectively enhance their capabilities, leading to improved overall performance. The research covers various aspects of multi-agent scenarios, including collective thinking, conversation dataset curation, sociological phenomenon exploration, and collaboration for efficiency. Collective thinking aims to boost problem-solving abilities by orchestrating discussions among multiple agents. Researchers like Wei et al. Wei et al. (2023) and Li et al. Li et al. (2023a) have created conversational datasets through role-playing methodologies. Sociological phenomenon investigations, such as Park et al. Park et al. (2023)’s work, involve creating virtual communities with rudimentary language interactions and limited cooperative endeavors. In contrast, Akata et al. Akata et al. (2023) scrutinized LLM cooperation through orchestrated repeated games. Collaboration for efficiency, proposed by Cai et al. Cai et al. (2023), introduces a model for cost reduction through large models as tool-makers and small models as tool-users. Zhang et al. Zhang et al. (2023) established a framework for verbal communication and collaboration, enhancing overall efficiency. However, Li et al. Li et al. (2023a) and Qian et al. Qian et al. (2023), presenting a multi-agent framework for software development, primarily relied on natural language conversations, not standardized software engineering documentation, and lacked advanced human process management expertise. Challenges in multi-agent cooperation include maintaining coherence, avoiding unproductive loops, and fostering beneficial interactions. Our approach emphasizes integrating advanced human processes, like code review in software maintenance, within multi-agent systems.
Appendix C Experimental Details
@@ C.1 Role Definition
C.2 Execute Time Across Languages
@@
Appendix D Comparative Analysis of QA-Checker AI System and Recursive Self-Improvement Systems
In this section, we will delve into the differences between QA-Checker and self-improvement systems Hong et al. (2023), and underscore the importance of the QA-Checker in role conversations.
D.1 Comparison Table
@@
| Feature/System | QA-Checker AI System | Recursive Self-Improvement System | ||||
@@
D.2 Differences and Implications
@@
D.3 Importance of QA-Checker in Role Conversations
@@ D.4 Conclusion
Appendix E Capabilities Analysis between CodeAgent and Other Methods
Compared to open-source baseline methods such as AutoGPT and autonomous agents such as ChatDev and MetaGPT, CodeAgent offers functions for code review tasks: consistency analysis, vulnerability analysis, and format analysis. As shown in Table 7, our CodeAgent encompasses a wide range of abilities to handle complex code review tasks efficiently. Incorporating the QA-Checker self-improved module can significantly improve the conversation generation between agents and contribute to the improvement of code review. Compared to COT, the difference and the advantages of CodeAgent with QA-Checker are shown in Section D.
Appendix F Dataset
@@
As shown in Zhou et al. (2023), our study incorporates three distinct datasets for evaluating the performance of CodeAgent: , , and . , compiled by Tufano et al. Tufano et al. (2021), derives from Gerrit and GitHub projects, excluding noisy or overly lengthy comments and review data with new tokens in revised code not present in the initial submission. , collected by Thongtanunam et al. Thongtanunam et al. (2022) from three Gerrit repositories, comprises only submitted and revised codes without review comments. Lastly, , gathered by Tufano et al. Tufano et al. (2022) from Java projects on GitHub, filters out noisy, non-English, and duplicate comments. These datasets are employed for Code Revision Before Review (CRB) and Code Revision After Review (CRA) tasks, with the exception of for CRA and Review Comment Generation (RCG) due to its lack of review comments.
@@
Our dataset, illustrated in Fig. 8, encapsulates a detailed analysis of consistency and format detection in software development, spanning various programming languages. It includes CA (consistency between commit and commit message (See Sec 2.1)) and FA (format consistency between commit and original (See Sec 2.1)) data, segmented into positive and negative samples based on the merged and closed status of pull requests. For example, in Python, the dataset comprises 254 merged and 35 closed negative CA samples, alongside 803 merged and 213 closed positive CA samples, with corresponding distributions for other languages like Java, Go, C++, and more. Similarly, the FA data follows this pattern of positive and negative samples across languages. Figure 7 graphically represents this data, highlighting the distribution and comparison of merged versus closed samples in both CA and FA categories for each language. This comprehensive dataset, covering over 3,545 commits and nearly 2,933 pull requests from more than 180 projects, was meticulously compiled using a custom crawler designed for GitHub API interactions, targeting post-April 2023 repositories to ensure up-to-date and diverse data for an in-depth analysis of current software development trends.
@@
Appendix G Key Factors Leading to Vulnerabilities
@@
| 1 | Insufficient Input Validation | Check for vulnerabilities like SQL injection, Cross-Site Scripting (XSS), and command injection in new or modified code, especially where user input is processed. |
@@
| 2 | Buffer Overflows | Particularly in lower-level languages, ensure that memory management is handled securely to prevent overflows. |
@@
| 3 | Authentication and Authorization Flaws | Evaluate any changes in authentication and authorization logic for potential weaknesses that could allow unauthorized access or privilege escalation. |
@@
| 4 | Sensitive Data Exposure | Assess handling and storage of sensitive information like passwords, private keys, or personal data to prevent exposure. |
@@
| 5 | Improper Error and Exception Handling | Ensure that errors and exceptions are handled appropriately without revealing sensitive information or causing service disruption. |
@@
| 6 | Vulnerabilities in Dependency Libraries or Components | Review updates or changes in third-party libraries or components for known vulnerabilities. |
@@
| 7 | Cross-Site Request Forgery (CSRF) | Verify that adequate protection mechanisms are in place against CSRF attacks. |
@@
| 8 | Unsafe Use of APIs | Check for the use of insecure encryption algorithms or other risky API practices. |
| 9 | Code Injection | Look for vulnerabilities related to dynamic code execution. |
@@
| 10 | Configuration Errors | Ensure that no insecure configurations or settings like open debug ports or default passwords have been introduced. |
@@
| 11 | Race Conditions | Analyze for potential data corruption or security issues arising from race conditions. |
@@
| 12 | Memory Leaks | Identify any changes that could potentially lead to memory leaks and resource exhaustion. |
@@
| 13 | Improper Resource Management | Check resource management, such as proper closure of file handles or database connections. |
@@
| 14 | Inadequate Security Configurations | Assess for any insecure default settings or unencrypted communications. |
@@
| 15 | Path Traversal and File Inclusion Vulnerabilities | Examine for risks that could allow unauthorized file access or execution. |
@@
| 16 | Unsafe Deserialization | Look for issues that could allow the execution of malicious code or tampering with application logic. |
| 17 | XML External Entity (XXE) Attacks | Check if XML processing is secure against XXE attacks. |
@@
| 18 | Inconsistent Error Handling | Review error messages to ensure they do not leak sensitive system details. |
@@
| 19 | Server-Side Request Forgery (SSRF) | Analyze for vulnerabilities that could be exploited to attack internal systems. |
@@
| 20 | Unsafe Redirects and Forwards | Check for vulnerabilities leading to phishing or redirection attacks. |
@@
| 21 | Use of Deprecated or Unsafe Functions and Commands | Identify usage of any such functions and commands in the code. |
@@
| 22 | Code Leakages and Hardcoded Sensitive Information | Look for hardcoded passwords, keys, or other sensitive data in the code. |
@@
| 23 | Unencrypted Communications | Verify that data transmissions are securely encrypted to prevent interception and tampering. |
@@
| 24 | Mobile Code Security Issues | For mobile applications, ensure proper handling of permission requests and secure data storage. |
@@
| 25 | Cloud Service Configuration Errors | Review any cloud-based configurations for potential data leaks or unauthorized access. |
Appendix H Data Leakage Statement
@@
As the new dataset introduced in Section F, the time of the collected dataset is after April 2023, avoiding data leakage while we evaluate CodeAgent on codeData dataset.
Appendix I Algorithmic Description of CodeAgent Pipeline with QA-Checker
@@
Appendix J Detailed Performance of CodeAgent in Various Languages on VA task
In our comprehensive analysis using CodeAgent, as detailed in Table 9, we observe a diverse landscape of confirmed vulnerabilities across different programming languages. The table categorizes these vulnerabilities into ‘merged’ and ‘closed’ statuses for languages such as Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. A significant finding is a markedly high number of ‘merged’ vulnerabilities in Python, potentially reflective of its extensive application or intrinsic complexities leading to security gaps. Conversely, languages like Go, Ruby, and C exhibit notably lower counts in both categories, perhaps indicating lesser engagement in complex applications or more robust security protocols. Table 9 that the ‘closed’ category consistently presents lower vulnerabilities than ‘merged’ across most languages, signifying effective resolution mechanisms. However, an exception is noted in C, where ‘closed’ counts surpass those of ‘merged’, possibly indicating either delayed vulnerability identification or efficient mitigation strategies. Remarkably, the Rateclose is generally observed to be higher than Ratemerge across the languages, exemplifying a significant reduction in vulnerabilities post-resolution. For example, Python demonstrates a Ratemerge of 14.00% against a higher Rateclose of 18.16%. This trend is consistent in most languages, emphasizing the importance of proactive vulnerability management. The Rateavg, representing the proportion of confirmed vulnerabilities against the total of both merged and closed items, further elucidates this point, with C++ showing the highest Rateavg at 16.49%. These insights not only underline the diverse vulnerability landscape across programming languages but also highlight the adeptness of CodeAgent in pinpointing and verifying vulnerabilities in these varied contexts.
@@
| Rateavg | 14.79% | 7.03% | 7.73% | 16.49% | 12.76% | 13.46% | 10.45% | 14.47% | 9.73% |
Appendix K More detailed experimental results on CA and FA tasks
@@
Appendix L Case Study
As shown in Table 10, we can easily localize the figure numbers of case studies for specific programming languages.
L.1 Performance on 9 languages
Figure No. | ||
| Python | 13 | ||
| Java | 14 | ||
| Go | 15 | ||
| C++ | 16 | ||
| JavaScript | 17 | ||
| C | 18 | ||
| C# | 19 | ||
| php | 20 | ||
| Ruby | 21 |
L.2 Difference of CodeAgent-3.5 and CodeAgent-4.0
CodeAgent-3.5 and CodeAgent-4.0 in this paper has no difference in general code review, however, CodeAgent-4.0 is more powerful in processing long input sequences and logic reasoning. As shown in Figure 22, we take one example of consistency detection between commit and commit message and find that CodeAgent-4.0 diffs from CodeAgent-3.5 in the detailed explanation. CodeAgent-3.5 output a report with 15k lines while CodeAgent-4.0 outputs a report with more than 17.7k lines. Detailed data is shown in https://zenodo.org/records/10607925.
Appendix M Ablation study
In this section, we evaluate the performance of different parts in CodeAgent in vulnerability analysis. CodeAgent is based on chain-of-thought (COT) and large language model (a.k.a. GPT). As shown in Section 4.1, CodeAgent outperforms baselines (a.k.a. CodeBERT, GPT-3.5, GPT-4.0) across 9 different languages. The performance mainly comes from the combination of COT and QA-Checker. Thus, we design an additional version called CodeAgent w/o, which means CodeAgent without QA-Checker. Then, we use CodeAgent w/o to do vulnerability analysis and compare with CodeAgent. We first discuss about the result of CodeAgent w/o and then discuss about comparison between CodeAgent and CodeAgent w/o.
Overview of Vulnerabilities in CodeAgent w/o
Table 11 presents the findings of CodeAgent w/o, a variant of the original CodeAgent, in identifying vulnerabilities across different programming languages. The table showcases the number of ‘merged’ and ‘closed’ vulnerabilities in languages such as Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. Notably, Python leads in the ‘merged’ category with a total of 1,057 cases, of which 140 are confirmed, yielding a Ratemerge of 13.25%. In contrast, languages like Go and Ruby show lower vulnerability counts in both ‘merged’ and ‘closed’ categories. The table also includes Rateclose and Rateavg, providing insights into the effectiveness of vulnerability management across these languages.
Detailed Comparison between CodeAgent and CodeAgent w/o
Comparing the findings in Table 11 with those in Table 9, we observe some notable differences in vulnerability detection by CodeAgent and CodeAgent w/o. While the overall trend of higher ‘merged’ vulnerabilities in Python and lower counts in Go and Ruby remains consistent, Table 11 shows a slight reduction in the Ratemerge for most languages, suggesting a more conservative confirmation approach in CodeAgent w/o. Similarly, Rateclose and Rateavg values in Table 11 generally indicate a lower proportion of confirmed vulnerabilities compared to Table 9, reflecting potentially different criteria or efficacy in vulnerability assessment. These variations highlight the impact of QA-Checker in CodeAgent.
@@
| Rateavg | 13.49% | 6.77% | 7.25% | 12.37% | 11.22% | 13.46% | 10.45% | 15.47% | 8.56% |
Appendix N Cost statement
As shown in Table 12, CodeAgent-4 has a higher query time and cost compared to CodeAgent-3.5 due to its increased complexity and capabilities. We acknowledge that the integration of AI models and multi-agent systems may introduce complexity and require specialized knowledge for implementation.
| Model | Query Time(min) | Cost in USD |
| CodeAgent-3.5 | 3 | 0.017 |
| CodeAgent-4 | 5 | 0.122 |
Appendix O Tool